OpenStudy (anonymous):

An ordinary (fair) die is a cube with the numbers through on the sides (represented by painted spots). Imagine that such a die is rolled twice in succession and that the face values of the two rolls are added together. This sum is recorded as the outcome of a single trial of a random experiment. Compute the probability of each of the following events: Event : The sum is greater than 5 . Event : The sum is divisible by 5.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

2 rolls are 36 possible outcomes; so it a/36 1,5; 5,1; 2,4; 4,2; 3,3 are all possible events that result in greater then 5 5/36

OpenStudy (amistre64):

the possible sums are 2 thru 12; so how many are divisible by 5? 5/5, 10/5 to make it 2/36

OpenStudy (anonymous):

@amistre64 I feel u r making a mistake P(sum greater than 5) = 26/36 = 13/18 and P(sum divisible by 5) = 7/36

OpenStudy (anonymous):

U can check by writing all possible 36 outcomes

OpenStudy (amistre64):

mistake? me? thats highly possible ;)

OpenStudy (amistre64):

1,5; 5,1; 2,4; 4,2; 3,3 ...... youre right, I stopped counting for some reason lol

OpenStudy (anonymous):

We all make mistakes every now n then, today it was your turn :)

OpenStudy (amistre64):

1,5 1,6 2,4 2,5 2,6 3,3 3,4 3,5 3,6 4,2 4,3 4,4 4,5 4,6 5,1 5,2 5,3 5,4 5,5 5,6 6,1 6,2 6,3 6,4 6,5 6,6 26/36

OpenStudy (anonymous):

Correct !!!!!

OpenStudy (anonymous):

By the way, are u a teacher??